OpenStudy (anonymous):

for alg.2 2 order pair to form a solution of the equation: 4x+3y=8 (0_,),(_0),(5,_)

jimthompson5910 (jim_thompson5910):

For (0,_), x = 0. So plug in x = 0 and solve for y 4x+3y=8 4(0)+3y=8 0+3y=8 3y=8 y=8/3 So when x = 0, y = 8/3, which means that the first ordered pair is (0, 8/3)

jimthompson5910 (jim_thompson5910):

Use this basic idea to find the other ordered pairs